ganeshie8 (ganeshie8):

I drive at an average speed of 30 miles per hour to the railroad station each morning and just catch my train. On a particular morning there was no traffic and at the halfway point I found I had averaged 60 miles per hour. How fast must I drive for the rest of the way to catch my train exactly without having to wait for the train ?

OpenStudy (reemii):

20?

ganeshie8 (ganeshie8):

Oh, that easy ?

ganeshie8 (ganeshie8):

Yeah that works ! how about this one : Next day, there was a lot of traffic and at the halfway point I found I had averaged only 15 miles per hour. How fast must I drive for the rest of the way to catch my train ?

OpenStudy (reemii):

\(\frac{D}{30} = T\) (T is time spent) \(T_1 + T_2 = T\) \(\frac{1/2}{60} + \frac{1/2}{S} = \frac{1}{30}\)

ganeshie8 (ganeshie8):

Nice!

OpenStudy (reemii):

Go fast!!

ganeshie8 (ganeshie8):

How fast

OpenStudy (reemii):

speed should be +infinity .